OpenStudy (steven513):

trey can clean a room in 40 minutes. his brother Marcus can clean the same room in 20 minutes. how long will it take them to clean the room if they work together?

OpenStudy (tkhunny):

You have "rooms per minute".. Change to "Minutes per room" and the world will be a clearer place.

OpenStudy (kasep10):

*

Directrix (directrix):

@steven513 Look at the attachment which shows the set-up for this type problem.

Directrix (directrix):

Trey does 1/40 of the job per minute Marcus does 1/20 of the job per minute x is the amount of time the two complete the job working togethr Following the pattern: 1/40 + 1/20 = 1/x 1/40 + 2/40 = 1/x 3/40 = 1/x Solve for x. Post what you get, okay? @steven513

OpenStudy (eliesaab):

Practice problems of this kind on http://www.saab.org/mathdrills/workcompletion.cgi